If someone says in 2 years they will be twice as old as they were 5years ago, how old are they?
Answer by ankor@dixie-net.com(22740)

Let a = his present age
:
"in 2 years they will be twice as old as they were 5years ago,"
a + 2 = 2(a-5)
a + 2 = 2a - 10
a - 2a = -10 - 2
-a = -12
a = 12 yrs is his present age
:
:
Check this in the given statement:
"in 2 years they will be twice as old as they were 5 yrs ago,"
12+2 = 2(12-5)
14 = 2(7)